OpenStudy (anonymous):

simplify the expression (x^3y^4 over x^-3y^7)^-2 A. y^3 over x^7 B. x^3 over y C. y^6 over x^10 D. x^-6 over y^-5 Please help

OpenStudy (anonymous):

\[\left(\frac{3^3y^4}{x^{-3}y^7}\right)^{-2}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you have a choice: you can multiply all the exponents by \(-2\) and then simplify the result, or you can simplify what is in the parentheses first and then multiply each exponent by \(-2\) makes no difference, which would you prefer?

OpenStudy (anonymous):

oh and i see i made a typo, it should be this: \[\left(\frac{x^3y^4}{x^{-3}y^7}\right)^{-2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets take care of what is inside the parentheses first

OpenStudy (anonymous):

subtracting the exponents (because you are dividing) and writing with positive exponents only gives \[\left(\frac{x^3y^4}{x^{-3}y^7}\right)^{-2}=\left(\frac{x^{3-(-3)}}{y^{7-4}}\right)^{-2}\]\[=\left(\frac{x^6}{y^3}\right)^{-2}\]

OpenStudy (anonymous):

now the exponent outside the parentheses has a minus sign, which we can get rid of by taking the reciprocal of what is inside the parentheses, that gives \[\left(\frac{y^3}{6^6}\right)^2\]

OpenStudy (anonymous):

well actually it gives \[\left(\frac{y^3}{x^6}\right)^2\]

OpenStudy (anonymous):

now square by multiplying each exponent by 2

OpenStudy (anonymous):

unfortunately none of your answer choices are right, so maybe there is either a typo in the question or in the answers